OpenStudy (anonymous):

Solve the following system by using graph paper or graphing technology. 2x + 2y = -6 3x - 2y = 11 What is the solution to the system? (-1, -7) (1, -4) (2, -1) (3, -2) WILL FAN AND GIVE MEDAL

OpenStudy (dan815):

add the 2 equations together

OpenStudy (dan815):

2x + 2y = -6 3x - 2y = 11 ------------- 5x+0y = 5

OpenStudy (dan815):

5x=5 x=1

OpenStudy (anonymous):

what would the ordered pair version of the answer be? I think thats what I have to find

OpenStudy (dan815):

sub value of x in and see what y is

OpenStudy (dan815):

2(1) + 2y = -6 or 3(1) - 2y = 11 solve for y

OpenStudy (anonymous):

so its C right

OpenStudy (dan815):

no

OpenStudy (anonymous):

i meant B

OpenStudy (dan815):

yes

OpenStudy (anonymous):

Thank you for your help